OpenStudy (anonymous):

x^4-4x^3+x^2+28x-56=0; √7 is a root Find the solution set. HELP

OpenStudy (anonymous):

factor it

OpenStudy (anonymous):

do u know how to factor ?

OpenStudy (anonymous):

?????? u there

OpenStudy (mathmate):

@bless Hints: If you know that \(\sqrt 7\) is a root, and seeing that 1. the coefficients are all integers, 2. constant term has 7 as a factor, you have a good reason to guess that -\(\sqrt 7\) is also a root. So try factor of (x+\(\sqrt 7\))(x-\(\sqrt 7\))=(x^2-7) by synthetic division. If it divides without remainder, you only have a quadratic to factor, if necessary.